CAN YOU PLEASE EXPLAIN THIS TO ME.
5 + 12 = 12 + 5

Answer by checkley71(8403)

5+12=12+5 OR 17=17
DO YOU HAVE A DOG (5) & A CAT (12)? IF YOU DO THEN YOU ALSO HAVE A CAT (12) & A DOG (5)!!!!!